# Deep Equilibrium Convolutional Sparse Coding for Hyperspectral Image Denoising

### v1

@@ IV-B1 ICVL Dataset
Figure 5: Row mean profiles of band 28 for the Nachal_08231214Nachal\_0823-1214 HSI from the ICVL dataset under the non-i.i.d. Gaussian noise with $\sigma\in[0,95]$. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
@@
[54] [55] [56] [17] [57] [25] [18] [30] [58] [31] [41] [59] (Ours)
@@
Figure 6: Denoising results on the the Houston 2018 HSI under the mixture noise. The false-color images are generated by combining bands 39, 20, and 4. (a) Clean. (b) Noisy. (c) BM4D [54]. (d) MTSNMF [55]. (e) LLRT [56]. (f) NGMeet [17]. (g) LRMR [57]. (h) E-3DTV [25]. (i) 3DlogTNN [18]. (j) SST [30]. (k) TRQ3D [58]. (l) SERT [31] (m) T3SC [41]. (n) MTSNN++ [59]. (o) DECSC.
@@ IV-B2 Houston 2018 HSI
Figure 7: Row mean profiles of band 39 for the Houston 2018 HSI under the mixture noise. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
Figure 8: Denoising results on the EO-1 HSI. The false-color images are generated by combining bands 97, 95, and 1. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
@@
Figure 9: Row mean profiles of band 159 for the EO-1 HSI. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
Figure 10: Denoising results of CapitalAirport HSI collected from the GF-5 satellite. The false-color images are generated by combining bands 153, 107, and 94. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
@@ IV-C1 Earth Observing-1 (EO-1) HSI
We selected an HSI captured by the EO-1 satellite, which covers a spectral range of 400 to 2500 nm. Following the experimental setup of Zhang et al. [57], a sub-image of size $200\times 400\times 166$ was used for testing. Fig. 8 compares the visual quality of restoration results produced by different methods. Most model-driven methods struggle to effectively remove stripe noise. E-3DTV, benefiting from its ability to capture correlations and differences across bands, produces relatively clear results even under severe noise. The differences between synthetic and real-world noise pose challenges for data-driven and hybrid-driven methods, leading to blurring in the results of methods like SST, TRQ3D, T3SC, and MTSNN++. In contrast, SERT and DECSC not only achieve effective denoising but also preserve well-defined edges and structural integrity in the reconstructions. The superior performance of SERT can be attributed to its rectangle self-attention mechanism, while DECSC’s advantage lies in the difference convolution’s ability to capture edge information, complemented by an attention mechanism that adaptively enhances critical features. Fig. 9 shows the row mean profiles of band 159 for the EO-1 HSI. Although clean ground truth is unavailable in this real-world scenario, the profiles still provide insight into the smoothness and consistency of the restored results. E-3DTV and DECSC achieve a favorable balance between smoothness and consistency with the value range of the noisy input, validating their robustness and effectiveness in practical scenarios.
@@ IV-C2 Gaofen-5 (GF-5) CapitalAirport HSI
Figure 11: Row mean profiles of band 154 for the CapitalAirport HSI collected from the GF-5 satellite. (a) Noisy. (b) BM4D [54]. (c) MTSNMF [55]. (d) LLRT [56]. (e) NGMeet [17]. (f) LRMR [57]. (g) E-3DTV [25]. (h) 3DlogTNN [18]. (i) SST [30]. (j) TRQ3D [58]. (k) SERT [31]. (l) T3SC [41]. (m) MTSNN++ [59]. (n) DECSC.
@@ IV-D2 Convergence Property
[54] [55] [56] [17] [57] [25] [18] [30] [58] [31] [41] [59] (Swin) (Mamba)

### v2

@@ IV-B1 ICVL Dataset
Figure 5: Row mean profiles of band 28 for the Nachal_08231214Nachal\_0823-1214 HSI from the ICVL dataset under the non-i.i.d. Gaussian noise with $\sigma\in[0,95]$. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
@@
[Maggioni2013BM4D] [Ye2015MTSNMF] [Chang2017LLRT] [He2022NGMeet] [Zhang2014LRMR] [Peng2020E-3DTV] [Zheng20203DlogTNN] [li2022spatialspectral] [Pang2022TRQ3DNet] [li2023spectral] [bodrito2021T3SC] [xiong2023multitask] (Ours)
@@
Figure 6: Denoising results on the the Houston 2018 HSI under the mixture noise. The false-color images are generated by combining bands 39, 20, and 4. (a) Clean. (b) Noisy. (c) BM4D [Maggioni2013BM4D]. (d) MTSNMF [Ye2015MTSNMF]. (e) LLRT [Chang2017LLRT]. (f) NGMeet [He2022NGMeet]. (g) LRMR [Zhang2014LRMR]. (h) E-3DTV [Peng2020E-3DTV]. (i) 3DlogTNN [Zheng20203DlogTNN]. (j) SST [li2022spatialspectral]. (k) TRQ3D [Pang2022TRQ3DNet]. (l) SERT [li2023spectral] (m) T3SC [bodrito2021T3SC]. (n) MTSNN++ [xiong2023multitask]. (o) DECSC.
@@ IV-B2 Houston 2018 HSI
Figure 7: Row mean profiles of band 39 for the Houston 2018 HSI under the mixture noise. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
Figure 8: Denoising results on the EO-1 HSI. The false-color images are generated by combining bands 97, 95, and 1. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
@@
Figure 9: Row mean profiles of band 159 for the EO-1 HSI. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
Figure 10: Denoising results of CapitalAirport HSI collected from the GF-5 satellite. The false-color images are generated by combining bands 153, 107, and 94. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
@@ IV-C1 Earth Observing-1 (EO-1) HSI
We selected an HSI captured by the EO-1 satellite, which covers a spectral range of 400 to 2500 nm. Following the experimental setup of Zhang et al. [Zhang2014LRMR], a sub-image of size $200\times 400\times 166$ was used for testing. Fig. 8 compares the visual quality of restoration results produced by different methods. Most model-driven methods struggle to effectively remove stripe noise. E-3DTV, benefiting from its ability to capture correlations and differences across bands, produces relatively clear results even under severe noise. The differences between synthetic and real-world noise pose challenges for data-driven and hybrid-driven methods, leading to blurring in the results of methods like SST, TRQ3D, T3SC, and MTSNN++. In contrast, SERT and DECSC not only achieve effective denoising but also preserve well-defined edges and structural integrity in the reconstructions. The superior performance of SERT can be attributed to its rectangle self-attention mechanism, while DECSC’s advantage lies in the difference convolution’s ability to capture edge information, complemented by an attention mechanism that adaptively enhances critical features. Fig. 9 shows the row mean profiles of band 159 for the EO-1 HSI. Although clean ground truth is unavailable in this real-world scenario, the profiles still provide insight into the smoothness and consistency of the restored results. E-3DTV and DECSC achieve a favorable balance between smoothness and consistency with the value range of the noisy input, validating their robustness and effectiveness in practical scenarios.
@@ IV-C2 Gaofen-5 (GF-5) CapitalAirport HSI
Figure 11: Row mean profiles of band 154 for the CapitalAirport HSI collected from the GF-5 satellite. (a) Noisy. (b) BM4D [Maggioni2013BM4D]. (c) MTSNMF [Ye2015MTSNMF]. (d) LLRT [Chang2017LLRT]. (e) NGMeet [He2022NGMeet]. (f) LRMR [Zhang2014LRMR]. (g) E-3DTV [Peng2020E-3DTV]. (h) 3DlogTNN [Zheng20203DlogTNN]. (i) SST [li2022spatialspectral]. (j) TRQ3D [Pang2022TRQ3DNet]. (k) SERT [li2023spectral]. (l) T3SC [bodrito2021T3SC]. (m) MTSNN++ [xiong2023multitask]. (n) DECSC.
@@ IV-D2 Convergence Property
[Maggioni2013BM4D] [Ye2015MTSNMF] [Chang2017LLRT] [He2022NGMeet] [Zhang2014LRMR] [Peng2020E-3DTV] [Zheng20203DlogTNN] [li2022spatialspectral] [Pang2022TRQ3DNet] [li2023spectral] [bodrito2021T3SC] [xiong2023multitask] (Swin) (Mamba)